Reduce into canonical form
4Uxx+5Uxy+Uyy+Ux+Uy=0
1
Expert's answer
2021-08-17T18:10:20-0400

"4Uxx+5Uxy+Uyy+Ux+Uy=0\\
or, 4r+5s+t+p+q=0\\
where R=4,S=5,T=1.\\
\text{therefore, characteristic equation is }R\lambda^2+S\lambda+T=0\\
4\lambda^2+5\lambda+1=0\\
\lambda=-1/4,-1\\
Then,\\
\text{the characteristics curve of given pde is the solution of these equation}\\
\frac{dy}{dx}+\lambda_1=0 --------(1)\\
and\\
\frac{dy}{dx}-\lambda_2=0---------(2)
\\
\text{from (1),we get}\\
c_1=y-(x/4) (=m)\\
\text{from (2),we get}\\
c_1=y-x (=n)\\
\text{Now, find p,q,r,s,t and substitute in the given pde.}\\
p=Ux=\frac{-1}{4}Um-Un\\
q=Uy=Um+Un\\
r=Uxx=\frac{1}{16}Umm+\frac{1}{2}Umn+Unn
\\
s=\frac{-1}{4}Umm-\frac{5}{4}Umn-Unn\\
t=Umm+2Umn+Unn\\
\text{substitute in the given pde, we get}\\
Umm=\frac{Um}{3}\\
\text{This is required canonical form.}"